# Language Models use Lookbacks to Track Beliefs

### v1

@@ 4.2 Uncovering the Binding Lookback Mechanism
Refer to captionCounterfactualOriginal
@@
Figure 4: Binding lookback payload and address: We intervened on both the high-level causal model and the LM running on the original story, modifying their variables and internal activations respectively, to match those from a counterfactual scenario. In the causal model intervention, we update the addresses (character and object OIs; Refer to caption and Refer to caption) and the payloads (state OIs; Refer to caption). This causes the binding lookback mechanism to attend to and retrieve the state OI corresponding to the alternate state token, which is then dereferenced by answer lookback to yield the alternate state token (e.g., beer instead of coffee). In the LM interchange intervention, modifying the residual stream at the state token results in identical outputs between layers 33 and 38. This confirms our hypothesis that both the address and payload information are represented in the residual stream of state tokens.
The Binding Lookback is the first operation applied to these OIs. The character and object OIs, serving as the source information, are each copied twice. One copy, referred to as the address, is placed in the residual stream of the state token (recalled token), alongside the state OI as the payload to transfer. The other copy, referred to as the pointer, is moved in the residual stream of the final token (lookback token). These pointer and address copies are then used to form the QK-circuit at the lookback token, which dereferences the state OI payload, transferring it from the state token to the final token. See Fig.2(a) (i) for a schematic of this lookback and see Fig.1 for the general mechanism.
Localizing the Address and Payload    In our first experiment, we localize the address copies of the character and object OIs and the state OI payload to the residual stream of the state token (recalled token, Fig. 2(a)). We sampled a counterfactual dataset where each example consists of an original input $\mathbf{o}$ with an answer that isn’t unknown and a counterfactual input $\mathbf{c}$ where the character, object, and state tokens are identical, except the ordering of the two sentences is swapped while the question remains unchanged, as illustrated in Fig. 4. The expected outcome predicted by our high-level causal model under intervention is the other state token from the original example, e.g., beer, because reversing the address and payload values without changing the pointer flips the output. In LM, the QK-circuit, formed using the pointer at the lookback token, attends to the other state token and retrieves its state OI as the payload.
@@
We perform an interchange intervention experiment layer-by-layer, where we replace the residual stream vectors at the first state token in the original run with that of the second state token in the counterfactual run and vice versa for the other state token. It is important to note that if the intervention targets state token values instead of their OIs, it should not produce the expected output. (This happens in the earlier layers.)
As shown in Fig. 4, the strongest alignment occurs between layers $33$ and $38$, supporting our hypothesis that the state token’s residual stream contains both the address information (character and object OIs) and the payload information (state OI). These components are subsequently used to form the appropriate QK and OV circuits of binding lookback.
@@
Localizing the Source Information    As shown in Fig. 2(a), the source information is copied as both the address and the pointer at different token positions. To localize the source information, we conduct intervention experiments with a dataset where the counterfactual example, c, swaps the order of the characters and objects as well as replaces the state tokens with entirely new ones while keeping the question the same as in o.
With this dataset, an interchange intervention on the high-level causal model that targets the source information will have downstream effects on both the address and the pointer, so no change in output occurs. However, if we additionally freeze the payloads and addresses, the causal model outputs the other state token, e.g., beer in Fig. 5, due to the mismatch between address and pointer.
In the LM, we interchange the residual streams of the character and object tokens while keeping the residual stream of the state token fixed. When the output of the intervened LM aligns with that of the intervened causal model, it indicates that the QK-circuit at the final token is attending to the alternate state token. As shown in Fig. 5, the second experiment reveals alignment between layers 20 and 34. This suggests that source information—specifically, the character and object OIs—is represented in their respective token residual streams within this layer range.
@@
Refer to captionCounterfactualOriginal
Figure 5: Source Information of Binding lookback: We run the causal model and the LM on an original story, then update their variables and activations, respectively, to the values they would take on for a counterfactual story with swapped characters and objects and new states. The interchange intervention on the high-level causal model swaps the sources of the binding lookback (character and object OIs; Refer to caption, Refer to caption) while freezing the addresses and payloads of the binding lookback (character, object, and state OIs; Refer to caption, Refer to caption, Refer to caption) and the answer lookback (state OI and token; Refer to caption, Refer to caption). By altering the sources, but freezing the addresses and payloads, only the pointer is changed so the binding lookback retrieves the other state OI which is dereferenced by the answer lookback to the other state token (e.g., beer instead of coffee). We perform the same interchange intervention on the LM and measure the agreement with the intervened causal model. Our results localize the source to the character and object token residual streams between layers 20 and 34.
We provide more experimental results in Appendix G where we show in Fig. 12 that freezing the residual stream of the state token is necessary. In sum, these results not only provide evidence for the presence of source information but also establish its transfer to the recalled and lookback tokens as addresses and pointers, respectively.
Localizing the Pointer Information    The pointer copies of the character and object OI are first formed at the character and object tokens in the question before being moved again to the final token for dereferencing (see Appendix H for experiments and more details).
4.3 Uncovering the Answer Lookback Mechanism
The LM answers the question using the Answer Lookback. The state OI of the correct answer serves as the source information, which is copied into two instances. One instance, the address copy of the state OI, is in the residual stream of the state token (recalled token) with the state token itself as the payload. The other instance, the pointer copy of the state OI, is transferred to the residual stream of the final token (lookback token) as the payload of the binding lookback. This pointer is then dereferenced, bringing the state token as the payload into the residual stream of the final token, which is predicted as the final output. See Fig. 2(a) (ii) for the answer lookback and Fig. 1 for the general mechanism.
Localizing the Pointer Information    We first localize the pointer of the answer lookback, which is the payload of the binding lookback. To do this, we conduct an interchange intervention experiment where the residual vectors at the final token position in the original run are replaced with those from the counterfactual run, one layer at a time. The counterfactual inputs have swapped objects and characters and randomly sampled states. If the answer pointer is targeted for intervention in the high-level causal model, the output is the other state in the original input, e.g., beer. As shown in Fig. 2(b), alignment begins at layer $34$, indicating that this layer contains pointer information, in low-rank subspace, which remains causally relevant until layer $52$.
Localizing the Payload    To determine where the model uses the state OI pointer to retrieve the state token, we use the same interchange intervention experiment. However, if the answer payload is targeted for intervention in the high-level causal model, the output is the correct state token from the counterfactual example, e.g., tea, rather than the state token from the original example, as illustrated in Fig. 2(b). The alignment occurs after layer $56$, indicating that the model retrieves the correct state token (payload) into the final token’s residual stream by $56$, where it is subsequently used to generate the final output.
@@ 5 Impact of Visibility Conditions on Belief Tracking Mechanism
In the previous section, we demonstrated how the LM uses ordering IDs and two lookback mechanisms to track the beliefs of characters that cannot observe each other. Now, we explore how the LM updates the beliefs of characters when provided with additional information that one of the characters (observing) can observe the actions of others (observed). We hypothesize that the LM employs another lookback mechanism, which we refer to as the Visibility Lookback, to incorporate information about the observed character.
@@
As illustrated in Fig. 6, we hypothesize that the LM first generates a Visibility ID at the residual stream of the visibility sentence, serving as the source information. The address copy of the visibility ID remains in the residual stream of the visibility sentence, while its pointer copy gets transferred to the residual streams of the subsequent tokens, which are the lookback tokens. Then LM forms a QK-circuit at the lookback tokens and dereferences the visibility ID pointer to bring forward the payload.
@@
Figure 6: Visbility Lookback When one character (the observing character) can see another (the observed character), the LM assigns a visibility ID (Refer to caption) to the visibility sentence (where this relation is defined). An address copy of this visibility ID remains in the visibility sentence’s residual stream. A pointer copy of the visibility ID is transferred to the subsequent tokens’ residual stream (lookback tokens). During processing, the model dereferences this pointer through a QK-circuit, bringing forward the payload ( Refer to caption ). Based on initial evidence, this payload contains the observed character’s OI(Refer to caption). Refer to Appendix I for more details. This mechanism allows the model to incorporate the observed character’s knowledge into the observing character’s belief state, enabling more complex belief reasoning.
@@
Although we were unable to determine the exact semantics of the payload in this lookback, we speculate that it encodes the observed character’s OI. We propose the existence of another lookback, where the story sentence associated with the observed character serves as the source, and its payload encodes information about the observed character. That payload information, encoding observed character’s OI, is then retrieved by the lookback tokens of the Visibility lookback, which contributes to the queried character’s enhanced awareness (see Appendix I for more details).
5.1 Uncovering the Visibility Lookback Mechanism
@@
Localizing the Source Information    To localize the source information, we conduct an interchange intervention experiment where the counterfactual is a different story with altered visibility information. In the original example, the first character cannot observe the second character’s actions, whereas in the counterfactual example, the first character can observe them (Fig. 7). The causal model outcome of this intervention is a change in the final output of the original run from “unknown” to the state token associated with the queried object. The interchange intervention is executed on visibility sentence tokens. As shown in Fig. 7 ( line), alignment occurs between layers $10$ and $23$, indicating that the visibility ID remains encoded in the visibility sentence until layer $23$, after which it is duplicated into address and pointer copies on visibility sentence and subsequent tokens respectively.
@@
Localizing the Payload    To localize the payload information, we use the same counterfactual dataset. However, instead of intervening on the source or recalled tokens, we intervene on the lookback tokens, specifically the question and answer tokens. As in the previous experiment, we replace the residual vectors of these tokens in the original run with those from the counterfactual run. As shown in Fig. 7 ( line), alignment occurs only after layer $31$, indicating that the information enhancing the queried character’s awareness is present in the lookback tokens only after this layer.
@@
Refer to captionCounterfactualOriginal
Figure 7: Visibility Lookback: We conduct three interchange intervention experiments to support the Visibility Lookback hypothesis: (1) Source Alignment: We align the source information (Refer to caption) by intervening on the visibility sentence—replacing it with its representation from a counterfactual run where the visibility sentence causes the queried character to become aware of the queried object’s contents. We observe that source information aligns between layers $10$ and $23$, after which it splits into separate address and pointer components. (2) Payload Alignment: To align the payload ( Refer to caption ), we intervene on all subsequent tokens and observe alignment only after layer $31$. (3) Address and Pointer Alignment: When intervening on both the address and pointer information (Refer to caption), we observe alignment across a broader range of layers, particularly between layers $24$ and $31$, because of the enhanced alignment between the address and pointer copies at the recalled and lookback tokens.
Localizing the Address and Pointer    The previous two experiments suggest the presence of a lookback mechanism, as there is no signal indicating that the source or payload has been formed between layers 24 and 31. We hypothesize that this lack of signal is due to a mismatch between the address and pointer information at the recalled and lookback tokens. Specifically, when intervening only on the recalled token after layer $25$, the pointer is not updated, whereas intervening only on the lookback tokens leaves the address unaltered, leading to the mismatch. To test this hypothesis, we conduct another intervention using the same counterfactual dataset, but this time, we intervene on the residual vectors of both the recalled and lookback tokens, i.e., the visibility sentence, as well as the question and answer tokens. As shown in Fig. 7 ( line), alignment occurs after layer $10$ and remains stable, supporting our hypothesis. This intervention replaces both the address and pointer copies of the visibility IDs, enabling the LM to form a QK-circuit and retrieve the payload.
@@ 6 Related Work
Theory of mind in LMs    A large body of work has focused on benchmarking different aspects of ToM through various tasks that attempt to assess LMs’ performance such as Le et al. (2019); Xu et al. (2024); Shapira et al. (2023); Jin et al. (2024); Wu et al. (2023); Kim et al. (2023b); Chan et al. (2024); Strachan et al. (2024b) and many more. In addition, there are various methods tailored to improve ToM ability in LMs through prompting (e.g., Sclar et al., 2023; Zhou et al., 2023; Wilf et al., 2024; Moghaddam & Honey, 2023; Hou et al., 2024).
Entity tracking in LMs    Entity tracking and variable binding are crucial abilities for LMs to exhibit not only coherent ToM capabilities, but also neurosymbolic reasoning. Many existing works have attempted to decipher this ability in LMs (Li et al., 2021; Davies et al., 2023; Kim & Schuster, 2023; Prakash et al., 2024; Feng & Steinhardt, 2023; Feng et al., 2024; Dai et al., 2024). Our work builds on their empirical insights and extends the current understanding of how LMs bind various entities defined in context.
Mechanistic interpretability of theory of mind    Only a few empirical studies explored the underlying mechanisms of ToM of LM (Zhu et al., 2024; Bortoletto et al., 2024) (Herrmann & Levinstein, 2024, is a notable theoretical paper). Those studies focus on probing techniques (Belinkov, 2022; Alain, 2016) to identify internal representations of beliefs and used steering techniques (Li et al., 2024; Rimsky et al., 2023) to improve LM performance by manipulating their activations. However, the mechanism by which LMs solve those tasks remains a black box, limiting our ability to understand, predict, and control LMs’ behaviors.
@@ 7 Conclusion
Through a series of desiderata-based patching experiments, we have mapped the mechanisms underlying the processing of partial knowledge and false beliefs in a set of simple stories. We are surprised by the pervasive appearance of a single recurring computational pattern: the lookback, which resembles a pointer dereference inside a transformer. The LMs use a combination of several lookbacks to reason about nontrivial visibility and belief states. Our improved understanding of these fundamental computations gives us optimism that it may be possible to fully reveal the algorithms underlying not only Theory of Mind, but also other forms of reasoning in LMs.

### v2

@@ 5.2 Step i: Binding Lookback – Linking Characters, Objects, and States
Refer to caption(a) Intervention Input Example(b) Layer-wise Intervention ResultsCounterfactualOriginal
@@
Figure 5: Binding lookback Address and Payload: The causal model predicts that swapping addresses (character and object OIs; Refer to caption and Refer to caption) and payloads (state OIs; Refer to caption) should cause the binding lookback mechanism to attend to the alternate state token and retrieve its state OI. This retrieved state OI is then dereferenced by the answer lookback, producing the corresponding token as the output (e.g., beer instead of coffee). The LM’s behavior matches this prediction when we perform interchange interventions on the state token across layers 33–38. These findings support our hypothesis that both address and payload information are encoded in the residual stream of state tokens.
@@
We perform the interchange intervention experiment layer-by-layer, where we replace the residual stream vector (or the identified subspace) of the first state token in the original run with that of the second state token in the counterfactual run and vice versa for the other state token. It is important to note that if the intervention targets state token values instead of their OIs, it should not produce the expected output. (This happens in the earlier layers.) As shown in Fig. 5b, the strongest alignment occurs between layers $33$ and $38$, supporting our hypothesis that the state token’s residual stream contains both the address (character and object OIs) and the payload information (state OI).
@@
Localizing the Source Reference Information    Next, we localize the source reference information, i.e., character and object OIs at their respective token residual stream. As illustrated in Fig. 6a, we conduct an intervention experiment with a dataset where the counterfactual example, c, swaps the order of the characters and objects as well as replaces the state tokens with entirely new ones while keeping the question the same as in o. Under this setup, an interchange intervention on the hypothesized causal model that targets the source reference should propagate changes through both the address and the pointer, leaving the final output unchanged. However, if we instead freeze the state token residual stream, which carries both the payload and the address, the causal model produces the alternate state token (e.g., beer in Fig. 6), as the pointer refers to the other state’s address.
In the LM, we interchange the residual streams of the character and object tokens layer-by-layer, while keeping the residual stream of the state token fixed. As shown in Fig. 6b, this experiment reveals alignment between layers $20$ and $34$, indicating that source reference is encoded in the residual streams of the character and object tokens within this layer range. Additional results are provided in Appendix G, where Fig. 13 shows that freezing the residual stream of the state token is necessary for this alignment to emerge. These findings support our hypothesis that source reference is present in the character and object tokens and is subsequently transferred to the recalled and lookback tokens.
@@
Refer to caption(a) Intervention Input Example(b) Layer-wise Intervention ResultsCounterfactualOriginal
Figure 6: Source Reference Information of Binding lookback: The causal model predicts that swapping the source reference information (character and object OIs; Refer to caption, Refer to caption), while freezing the addresses and payloads of the binding lookback, should cause the binding lookback mechanism to attend to the alternate state token and retrieve its state OI, which would generate alternate state token as the final output via the answer lookback (e.g., beer instead of coffee). The LM’s behavior matches this prediction when we perform interchange interventions at the character and object tokens across layers 20-34. These results support our hypothesis that source reference information is encoded in the residual stream of character and object tokens.
Localizing the Pointer Information    Finally, we localize the pointer copies of the character and object OIs to their corresponding tokens in the question and to the final token. See Appendices G &  H for details of the experiments and results.
In summary, belief tracking begins in layers 20–34, where character and object OIs are encoded in their respective token representations. These OIs are transferred to the corresponding state tokens in layers 33–38. When a question is asked, pointer copies of the relevant character and object OIs are moved to the final token by layer 34, where they are dereferenced to retrieve the correct state OI. At the final token, this state OI is represented across layers 34–52, and between layers 52–56, it is dereferenced to fetch the answer from the correct state token, producing the final output.
@@ 6 Impact of Visibility Conditions on Belief Tracking Mechanism
So far, we have demonstrated how the LM uses ordering IDs and two lookback mechanisms to track the beliefs of characters that cannot observe each other. Now, we explore how the LM updates the beliefs of characters when one character (observing) can observe the actions of the other (observed).
Hypothesized Visibility Lookback Mechanism
@@
We hypothesize that the LM uses an additional lookback mechanism, which we call the Visibility Lookback, to integrate information about the observed character when it is explicitly stated that one character can see another’s action. As illustrated in Fig. 7, we hypothesize that the LM first generates a Visibility ID ([Uncaptioned image]) at the residual stream of the visibility sentence, serving as the source reference information. The address copy of the visibility ID remains in the residual stream of the visibility sentence, while its pointer copy gets transferred to the residual stream of the subsequent tokens (lookback tokens). Then LM forms a QK-circuit at the lookback tokens and dereferences the visibility ID pointer to retrieve the payload.
@@
Figure 7: Visibility Lookback When one (observing) character can see another (observed) character, the LM assigns a visibility ID (Refer to caption) to the visibility sentence where this relation is defined. An address copy of this visibility ID remains in the visibility sentence’s residual stream. A pointer copy of the visibility ID is transferred to the subsequent tokens’ residual stream. The LM dereferences this pointer through a QK-circuit, bringing forward the payload ( Refer to caption ), when processing subsequent tokens. Based on initial evidence, this payload contains the observed character’s OI(Refer to caption). See Appendix I for details. This mechanism allows the model to incorporate the observed character’s knowledge into the observing character’s belief state, enabling more complex belief reasoning.
@@
Although our two-character setting is unable to discern the exact semantics of the payload in the visibility lookback, our observations are consistent with the payload encoding the observed character’s OI. Our initial observations suggest another lookback where the story sentence associated with the observed character serves as the source reference, and its payload encodes information about the observed character. The observed character’s OI appears to be retrieved by the lookback tokens of the Visibility lookback, with causal effects on the queried character’s awareness (see App. I for details).
6.1 Verifying the Hypothesized Visibility Lookback
@@
Localizing the Source Reference    In this experiment, we localize the Visibility ID ([Uncaptioned image]), i.e., the source reference of the Visibility lookback. We conduct an interchange intervention experiment where the counterfactual is a different story in which the characters’ visibility is flipped from unobserved to observed (Fig. 8a), and we look for an output change from “unknown” to the answer that would be observed. We intervene on the representation of all the visibility sentence tokens. As shown in Fig. 8b (blue line), causal effects appear between layers $10$ and $23$, indicating that the visibility ID remains encoded in the visibility sentence until layer $23$, after which it is split into address and pointer copies that must be connected by dereference to have an effect. This pattern supports our hypothesis that the LM generates a reference to the Visibility ID.
@@
Localizing the Payload    Next, we localize the payload ( [Uncaptioned image] ) information using the same counterfactual dataset. However, instead of intervening on the recalled tokens, we intervene on the lookback tokens, specifically the question and answer tokens. As in the previous experiment, we replace the residual vectors of these tokens in the original run with those from the counterfactual run. As shown in Fig. 8b (red line), alignment occurs after layer $31$, indicating that the information causing the queried character’s awareness is present in the lookback tokens after this layer.
@@
Refer to caption(a) Intervention Input Example(b) Layer-wise Intervention ResultsCounterfactualOriginal
Figure 8: Visibility Lookback: We conduct three interchange intervention experiments to support the Visibility Lookback hypothesis: (1) Source Alignment: We align the source reference (Refer to caption) by intervening on the visibility sentence, replacing it with its representation from a counterfactual run where the visibility sentence causes the queried character to become aware of the queried object’s contents. We observe that source reference information aligns between layers $10$ and $23$, after which it splits into separate address and pointer components. (2) Payload Alignment: To align the payload ( Refer to caption ), we intervene on all subsequent tokens and observe alignment only after layer $31$. (3) Address and Pointer Alignment: When intervening on both the address and pointer information (Refer to caption), we observe alignment across a broader range of layers, particularly between layers $24$ and $31$, because of the enhanced alignment between the address and pointer copies at the recalled and lookback tokens.
Localizing Address and Pointer    The previous two experiments indicate the absence of both the source and payload information between layers $24$ and $31$. We hypothesize that this lack of signal is due to a mismatch between the address and pointer information that inhibits a lookback dereference. Specifically, when intervening only on the recalled token after layer $25$, the pointer is not updated, whereas intervening only on the lookback tokens leaves the address unaltered, a mismatch in either case. To test this hypothesis, we conduct another intervention using the same counterfactual dataset, but this time, we intervene on the residual vectors of both the recalled and lookback tokens, i.e., the visibility sentence, as well as the question and answer tokens. As shown in Fig. 8b (green line), alignment occurs after layer $10$ and remains stable, providing evidence that a lookback now occurs between layers 24 and 31. This intervention replaces both the address and pointer copies of the visibility IDs, enabling the LM to form a QK-circuit and resolve the visibility lookback.
@@ 7 Related Work
Theory of mind in LMs    Theory of mind in LMs has been widely benchmarked (Le et al., 2019; Shapira et al., 2023; Wu et al., 2023; Kim et al., 2023; Xu et al., 2024; Jin et al., 2024; Chan et al., 2024; Strachan et al., 2024b). However, these benchmarks lack adequate counterfactuals for the binding manipulations we need, so we made CausalToM (Section B).
Entity tracking in LMs    Entity tracking and variable binding are crucial abilities for LMs to exhibit not only coherent ToM capabilities, but also neurosymbolic reasoning. Many existing works have attempted to decipher this ability in LMs (Li et al., 2021; Davies et al., 2023; Feng & Steinhardt, 2023; Kim & Schuster, 2023; Prakash et al., 2024; Feng et al., 2024; Dai et al., 2024; Wu et al., 2025). Our work builds on their empirical insights and extends the current understanding of how LMs bind various entities defined in context.
Mechanistic interpretability of theory of mind    Few studies explored the underlying mechanisms of ToM of LM (Zhu et al., 2024; Bortoletto et al., 2024; Herrmann & Levinstein, 2024). Those studies use probing (Alain, 2016; Belinkov, 2022) to identify internal representations of beliefs and steering (Rimsky et al., 2023; Li et al., 2024) to control LMs by manipulating their activations. However, the mechanism by which LMs solve those tasks remains a black box, limiting our ability to understand, predict, and control LMs’ behaviors.
@@ 8 Conclusion
Through a series of interchange intervention experiments, we have mapped the end-to-end underlying mechanism responsible for the processing of partial knowledge and false beliefs in a set of simple stories. We are surprised by the pervasive appearance of a single recurring computational pattern: the lookback, which resembles a pointer dereference inside a transformer. The LMs use a combination of several lookbacks to reason about nontrivial belief states. Our improved understanding of these fundamental computations gives us optimism that it may be possible to fully reveal the algorithms underlying not only the theory of mind, but also other capabilities in LMs.
9 Ethics Statement
This work involves experiments conducted exclusively on synthetic text generated by the authors. No human subjects, personal data, or sensitive user information were collected or analyzed. The models we analyze are publicly released LLMs (Llama-3-70B-Instruct and Llama-3.1-405B-Instruct). All interventions and analyses were performed locally without querying proprietary APIs. We acknowledge that research on ToM in LMs may be misinterpreted as suggesting human-like cognition or intentionality. We explicitly caution that our findings demonstrate internal computational mechanisms, not conscious reasoning, and should not be construed as evidence of sentience or moral agency in LMs. Our causal intervention techniques reveal latent structures within pretrained models but do not modify the underlying weights or enable novel capabilities. Nevertheless, reverse-engineering latent belief representations could, in principle, be misused for behavioral steering or manipulation. To mitigate this, we provide our code strictly for transparency and further scientific audit, rather than for application in deployed systems.
10 Reproducibility Statement
To facilitate reproducibility, we release the full CausalToM dataset, including all story templates and the code used to generate the various story instances that serve as counterfactual variants in our experiments. The repository, that can be accessed at https://belief.baulab.info, contains all scripts required to construct the dataset, extract activations, perform interchange interventions, and compute causal mediation metrics, along with the hyperparameters and random seeds used for subspace identification via DCM. All experiments were conducted using publicly available open-weight models (Llama-3-70B-Instruct and Llama-3.1-405B-Instruct). The interchange intervention experiments can be reproduced either by hosting these models locally or by using remotely hosted instances via NDIF (Fiotto-Kaufman et al., 2025), if sufficient local compute is not available.
11 The Use of Large Language Models
We used LLMs as a writing assistant to correct grammatical and typographical errors; beyond this, they did not contribute to any stage of the research.